OpenStudy (anonymous):

Express the given quantity as a single logarithm. ln(6) + 4 ln(4)

OpenStudy (anonymous):

note that a log x = log x^a and log a + log b = log ab

OpenStudy (anonymous):

\[\ln (6)+4\ln(4)=\ln(6)+\ln(4^4)=\ln(6.4^4)=\ln(1536)\]

OpenStudy (anonymous):

\[=\ln(6)+\ln4^{4}=\ln(6*4^{4})\]